OpenStudy (yanasidlinskiy):

A standard deck of cards contains 52 cards. One card is selected from the deck. ​(a) Compute the probability of randomly selecting a diamond or spade. ​(b) Compute the probability of randomly selecting a diamond or spade or club. ​(c) Compute the probability of randomly selecting a queen or club.

OpenStudy (yanasidlinskiy):

Would it be (13/52)(13/52)?? Or (1/52)(1/52)

OpenStudy (raffle_snaffle):

^that is part a

OpenStudy (yanasidlinskiy):

Wait, how did you get the 48 if I'm only selecting a diamond or a spade though?

OpenStudy (raffle_snaffle):

there are 4 diamonds of spaded in a deck of 52 cards

OpenStudy (zarkon):

P(diamond or spade)=P(diamond)+P(spade)

OpenStudy (yanasidlinskiy):

Ahh ok.

OpenStudy (yanasidlinskiy):

Wait, I'm confused. Like there's 10 spades and diamonds without counting the jacks and queens.

OpenStudy (zarkon):

there are 13 spades and 13 diamonds

OpenStudy (yanasidlinskiy):

Yes, but I wasn't counting all of the spades and diamonds that are included with the Jacks and Queens

OpenStudy (zarkon):

why?

OpenStudy (yanasidlinskiy):

Well aren't I supposed to be counting ONLY spades and diamonds?

OpenStudy (zarkon):

queen of spades is a spade

OpenStudy (yanasidlinskiy):

Well okay. I've had that pulled up already but thanks.

OpenStudy (zarkon):

there are 13 hearts, 13 clubs, 13 diamonds and 13 spades

OpenStudy (yanasidlinskiy):

yes, I do know that. And randomly selecting a diamond or spade would be 4?

OpenStudy (zarkon):

4?

OpenStudy (yanasidlinskiy):

*2

OpenStudy (zarkon):

I still don't know what you are doing

OpenStudy (zarkon):

half the deck is a spade or diamond

OpenStudy (zarkon):

that is 26/52=1/2

OpenStudy (yanasidlinskiy):

Ok, that makes total sense now. Better off than telling me random numbers and getting me all confused. So for part A it'd be: 1/2 right? since spades and diamonds is half the deck right?

OpenStudy (zarkon):

yes

OpenStudy (yanasidlinskiy):

Ok. hang on. I'm gonna try to figure out B on my own.

OpenStudy (yanasidlinskiy):

Before I go on it'd be 39/52 right? Since there's 13 of each.. If not, I'd like an explanation on what I'm supposed to do.

OpenStudy (zarkon):

that is correct 39/52=3/4

OpenStudy (yanasidlinskiy):

So, .75. And for C selecting a queen or club would be: (4/52)(12/52) Correct?

OpenStudy (yanasidlinskiy):

Wait, I have a question.

OpenStudy (zarkon):

P(A or B)=P(A)+P(B)-P(A and B)

OpenStudy (yanasidlinskiy):

Would I still be counting the spade that is within the queen?

OpenStudy (yanasidlinskiy):

Cuz there's 4 queens and 13 spades...

OpenStudy (zarkon):

there are...and one of those is the queen of spades

OpenStudy (yanasidlinskiy):

So would be counting it then right?

OpenStudy (zarkon):

we would ...just don't count it twice

OpenStudy (yanasidlinskiy):

Ahh, ok. Gotcha. So setting it up.. I'd have: (4/52)(12/52) right?

OpenStudy (zarkon):

no...you do not multiply

OpenStudy (yanasidlinskiy):

Subtract?

OpenStudy (yanasidlinskiy):

Or add, I don't know.

OpenStudy (zarkon):

well it's one of those

OpenStudy (yanasidlinskiy):

Well, you have me this formula sooo.... I'm guessing subtract then.. P(A or B)=P(A)+P(B)-P(A and B)

OpenStudy (zarkon):

P(queen or club)=P(queen)+P(club)-P(queen and club)

OpenStudy (yanasidlinskiy):

OMG, this is so confusing. I don't know.

OpenStudy (zarkon):

P(queen)=?

OpenStudy (yanasidlinskiy):

Queen= 4/52 Club = 12/52

OpenStudy (zarkon):

P(club)=13/52

OpenStudy (zarkon):

P(queen and club)=?

OpenStudy (yanasidlinskiy):

I thought we weren't counting the spades twice.

OpenStudy (zarkon):

we count each unrestricted and we subtract any we double counted. That is what the formula does

OpenStudy (yanasidlinskiy):

OHh ok. that makes sense so 17/52?

OpenStudy (zarkon):

almost P(Q or C)=P(Q)+P(C)-P(Q and C) \[=\frac{4}{52}+\frac{13}{52}-\frac{1}{52}=\frac{16}{52}\]

OpenStudy (yanasidlinskiy):

Oh, could't I have subtracted that in the beginning and still come out to the same answer?

OpenStudy (zarkon):

yes...but you can't write P(club)=12/52 since the real answer is P(club)=13/52

OpenStudy (yanasidlinskiy):

Oh, ok. I see where that's coming from. Anyways thank you soo much. I doo appreciate it.